MICROWAVE SPECTRA AND GEOMETRIES OF PtCO AND PdCO

COREY J. EVANS, NICHOLAS R. WALKER, JOSEPH K. H. LUI and \underline{MICHAEL C. L. GERRY}, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, B. C., Canada, V6T 1Z1.

Pure rotational spectra of several isotopomers of PtCO and PdCO have been observed in the frequency range 6-20 GHz, using a cavity pulsed jet Fourier transform microwave spectrometer. The samples were prepared by laser ablation of Pt or Pd metal in the presence of 1-1.5% CO contained in the Ar backing gas of the jet. Assignments were confirmed from the isotopic dependence of the spectra, including, where appropriate, spin-rotation and nuclear quadrupole hyperfine patterns. Structures of both molecules have been determined, and will be discussed in the context of those of other metal carbonyls.